OpenStudy (anonymous):

convergence of (-1)^k x lnk/k

OpenStudy (zarkon):

\[\left\{\frac{(-1)^kx\ln(k)}{k}\right\}_{k=1}^{\infty}\] it is conditionally convergent for all x not equal to zero and absolutely convergent when x=0